how do you solve this linear system by using substitution
x=6-4y
2x-3y=1

Answer by jim_thompson5910(35256)

Hint:

2x-3y=1

2(6-4y)-3y=1 ... plug in x=6-4y

12 - 8y - 3y = 1

12 - 11y = 1

Keep going to solve for y. Then use this value of y to find the value of x.